# supplemental-material

### v1

@@ Abstract
On the other hand, it is known that axions can serve as good dark matter (DM) candidate Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . This motivates us to investigate how the axion-mediated forces are affected in the presence of axion DM as background. A similar effect has been studied in the context of neutrino forces in the presence of cosmic neutrino background Horowitz:1993kw ; Ferrer:1998ju ; Ferrer:1999ad , and has recently been extended to various neutrino backgrounds Ghosh:2022nzo ; Arvanitaki:2022oby ; Arvanitaki:2023fij ; Ghosh:2024qai as well as dark forces induced by the exchanging light dark sector particles Barbosa:2024pkl ; VanTilburg:2024xib . An inspiring interpretation of the background-related force was introduced in VanTilburg:2024xib , termed as the “wake force”, offering insights from both classical and quantum perspectives. The presence of a background field drastically impacts processes at the loop level rather than the tree level. For instance, the anomalous magnetic moment of Standard Model (SM) particles receives a large correction from a DM background field Evans:2023uxh ; Arza:2023wou ; Evans:2024dty . This is basically because when doing a loop integral, the number density of background particles enters, which can dominate over the vacuum counterpart. Consequently, axion-mediated forces are especially vulnerable to the background correction, since there is no tree-level diagrams involved.
To reiterate, when the number density of background particles becomes sufficiently large, zero-temperature field theory is no longer valid. The axion DM mass is in general very light, ranging from $10^{-22}\,$eV to $O(1)\,$eV (see e.g., Refs. Marsh:2015xka ; Lin:2019uvt ), implying that the number density of axion DM is enormously large with the DM energy density known to be $\sim 0.3{\rm~{}GeV}/{\rm cm}^{3}$. The axion propagator must be modified with such a background present. In the vacuum-to-vacuum case, the exception value of the normal ordered product of annihilation and creation operators is a delta function, $\langle 0|a_{\boldsymbol{p}}a^{\dagger}_{\boldsymbol{k}}|0\rangle=(2\pi)^{3}\delta^{3}(\boldsymbol{p}-\boldsymbol{k})$. However, for an axion emitted from a background axion “sea” $|n\rangle$, propagating a distance, and then reabsorbed into the background, the expectation becomes
@@
We show that the background effect changes the scaling of the potential for the axion force from $1/r^{3}$ to $1/r$. Also, the magnitude of the potential is greatly enhanced, which becomes proportional to the number density of axion DM particles. Consequently, for a wide range of axion mass, the constraints on the axion parameter $c_{GG}/f$ is improved by several orders of magnitude based on the fifth-force experiments such as the Casimir-less and torsion balance tests.
@@
Here, $\tau\equiv m_{a}^{2}/m_{\pi}^{2}$, the parameters $c_{u,d}$ and $c_{GG}$ are the coupling of axion with quarks and gluon in UV theory with coefficients $g_{0}=0.440(44)$ and $g_{A}=1.254(16)(30)$ Liang:2018pis ; Bauer:2021mvw . Eq. (3) preserves the axion shift symmetry except for the axion mass term. The shift-symmetry breaking effect will be reflected in higher orders. At the quadratic order of $f^{-2}$, the axion-nucleon interactions can be described by the dimensional six operators Gasser:1987rb ; Bauer:2023czj ; Bauer:2024hfv
@@
All five types of diagrams depicting the exchange of an axion pair between two fermions $N_{1}$ and $N_{2}$ are listed in Fig. 1, which contribute to the axion force. The three-point vertex represents the linear coupling $-im_{N}g_{N}/f$ as shown in the first term of Eq. (7). There are two types of four-point vertices, $m_{N}g_{N}^{2}/2f^{2}$ and $c_{N}g_{N}^{2}/f^{2}$, corresponding to the second and the third terms in Eq. (7), respectively. We call the second term the shift-symmetry-invariant (SI) operator while the third term the shift-symmetry-breaking (SB) operator. One should not be confused by the terminology. The first two terms in Eq. (7) are the leading two orders in the Taylor expansion, which roots in the partial-derivative coupling term in Eq. (3) that respects the axion shift symmetry. On the other hand, the third term arises due to the shift-symmetry breaking effect, as shown in Eq. (6).
@@
Assuming an isotropic distribution $n(\boldsymbol{k})$ of DM background, we can integrate out the angular part of $\boldsymbol{k}$. Eventually, we get the following expression of the background-enhanced potential $V(r)$ for the axion force,
@@
Based on these experiments, we can finally set constraints on the axion parameter $c_{GG}/f$, and the results are shown as solid lines in Fig. 2. For comparison, the constraints by considering only the vacuum contribution to the force are shown in dash-dotted lines. The blue lines are for the CL experiment, while the red lines are for the torsion-balance inverse-square law tests. As we can see, with the background effect included, the constraints are several orders of magnitude more stringent than in the vacuum-only case. Additionally, as the axion mass decreases, the constraints become increasingly stringent, scaling as $m_{a}^{-1/2}$. The enhancement of constraints comes from two facts. First, the background effect modifies the scaling of the potential from $1/r^{3}$ to $1/r$. Second, the effect is proportional to the number density of axion DM and is inversely proportional to the axion mass. The dashed extension lines attached to the solid lines represent the parameter region where the DM de Broglie wavelength is comparable with the separation between the test and source masses in the two experiments. As $m_{a}$ becomes larger and thus the de Broglie wavelength becomes shorter, the constraints are expected to smoothly transition back to the vacuum case.
Discussion and conclusion– In this work, we have investigated the impact of the axion DM background on the spin-independent axion forces between nucleons. This is particularly important for the axion case, as the background effect modifies only the loop-level results, and the axion-mediated forces themselves are loop-level phenomena. By systematically calculating the relevant Feynman diagrams, we have demonstrated that the axion force is significantly amplified in proportion to the number density of axion DM particles. Furthermore, the scaling of the potential for this force is modified from $1/r^{3}$, as in the vacuum-only case, to $1/r$ for distance $r$ smaller than the DM de Broglie wavelength. Consequently, across a broad range of axion masses, constraints on the axion parameter $c_{GG}/f$ have been significantly enhanced by existing results from fifth-force experiments, including Casimir-less and torsion balance tests. This underscores the fact that these types of experiments are more effective than previously believed in searching for the hypothetical particle, the axion.
An interesting future direction would involve investigating how the background effect diminishes when the typical length scale of experiments approaches or exceeds the DM de Broglie wavelength. Additionally, it should be noted that for simplicity, our calculations assume an isotropic distribution of DM. A more meticulous analysis may reveal daily and/or annual modulations of the enhanced force.
@@ Acknowledgements
This work is partly supported by the National Research Foundation of Korea (NRF) under Grant No. RS-2024-00405629.

### v2

@@ Abstract
On the other hand, it is known that axions can serve as good dark matter (DM) candidate Preskill:1982cy ; Abbott:1982af ; Dine:1982ah . This motivates us to investigate how the axion-mediated forces are affected in the presence of axion DM as background. A similar effect has been studied in the context of neutrino forces in the presence of cosmic neutrino background Horowitz:1993kw ; Ferrer:1998ju ; Ferrer:1999ad , and has recently been extended to various neutrino backgrounds Ghosh:2022nzo ; Arvanitaki:2022oby ; Arvanitaki:2023fij ; Ghosh:2024qai as well as dark forces induced by the exchanging light dark sector particles Barbosa:2024pkl ; VanTilburg:2024xib . An inspiring interpretation of the background-related force was introduced in VanTilburg:2024xib , termed as the “wake force”, offering insights from both classical and quantum perspectives. The presence of a background field drastically impacts processes at the loop level rather than the tree level. For instance, the anomalous magnetic moment of Standard Model (SM) particles receives a large correction from a DM background field111 A classical description is proposed in Zhou:2025wax , drawing an analogy to the ponderomotive force in optics. The author argues that the effect becomes significant only when the DM mass exceeds the characteristic frequency $\omega_{0}$ of the $g-2$ experimental. Evans:2023uxh ; Arza:2023wou ; Evans:2024dty . This is basically because when doing a loop integral, the number density of background particles enters, which can dominate over the vacuum counterpart. Consequently, axion-mediated forces are especially vulnerable to the background correction, since there is no tree-level diagrams involved. See also other types of DM background effects Fukuda:2021drn ; Day:2023mkb ; Li:2024bbe ; Du:2024tin ; Yin:2023jjj ; Alonso-Alvarez:2019ssa .
To reiterate, when the number density of background particles becomes sufficiently large, zero-temperature field theory is no longer valid. The axion DM mass is in general very light, ranging from $10^{-22}\,$eV to $O(1)\,$eV (see e.g., Refs. Marsh:2015xka ; Lin:2019uvt ), implying that the number density of axion DM is enormously large with the DM energy density known to be $\sim 0.4{\rm~{}GeV}/{\rm cm}^{3}$. The axion propagator must be modified with such a background present. In the vacuum-to-vacuum case, the exception value of the normal ordered product of annihilation and creation operators is a delta function, $\langle 0|a_{\boldsymbol{p}}a^{\dagger}_{\boldsymbol{k}}|0\rangle=(2\pi)^{3}\delta^{3}(\boldsymbol{p}-\boldsymbol{k})$. However, for an axion emitted from a background axion “sea” $|n\rangle$, propagating a distance, and then reabsorbed into the background, the expectation becomes
@@
In this paper, we show that the background effect changes the scaling of the potential for the axion force from $1/r^{3}$ to $1/r$. Also, the magnitude of the potential is greatly enhanced, which becomes proportional to the number density of axion DM particles. Consequently, for a wide range of axion mass, the constraints on the axion parameter $c_{GG}/f$ is improved by several orders of magnitude based on the fifth-force experiments such as the Casimir-less and torsion balance tests.
@@
Here, $\tau\equiv m_{a}^{2}/m_{\pi}^{2}$, the parameters $c_{u,d}$ and $c_{GG}$ are the coupling of axion with quarks and gluon in UV theory with coefficients $g_{0}=0.440(44)$ and $g_{A}=1.254(16)(30)$ Liang:2018pis ; Bauer:2021mvw . Eq. (3) preserves the axion shift symmetry except for the axion mass term. The shift-symmetry breaking effect will be reflected in higher orders. At the quadratic order of $f^{-2}$, the shift symmetry breaking axion-nucleon interactions can be described by the dimensional six operators Gasser:1987rb ; Bauer:2023czj ; Bauer:2024hfv
@@
All five types of diagrams depicting the exchange of an axion pair between two fermions $N_{1}$ and $N_{2}$, which contribute to the axion force, are listed in Fig. 1. The three-point vertex represents the linear coupling $-im_{N}g_{N}/f$ as shown in the first term of Eq. (7). There are two types of four-point vertices, $m_{N}g_{N}^{2}/2f^{2}$ and $c_{N}g_{N}^{2}/f^{2}$, corresponding to the second and the third terms in Eq. (7), respectively. We call the second term the shift-symmetry-invariant (SI) operator while the third term the shift-symmetry-breaking (SB) operator. One should not be confused by the terminology. The first two terms in Eq. (7) are the leading two orders in the Taylor expansion, which roots in the partial-derivative coupling term in Eq. (3) that respects the axion shift symmetry. On the other hand, the third term arises due to the shift-symmetry breaking effect, as shown in Eq. (5).
@@
Assuming an isotropic distribution $n(\boldsymbol{k})$ of DM background, we can integrate out the angular part of $\boldsymbol{k}$. Fourier transforming back to the position space, we eventually get the following expression of the background-enhanced potential $V(r)$ for the axion force,
@@
Based on these experiments, we can finally set constraints on the axion parameter $c_{GG}/f$, and the results are shown as solid lines in Fig. 2. For comparison, the constraints by considering only the vacuum contribution to the force are shown in dash-dotted lines. The blue lines are for the CL experiment, while the red lines are for the torsion-balance inverse-square law tests. As we can see, with the background effect included, the constraints are several orders of magnitude more stringent than in the vacuum-only case. Additionally, as the axion mass decreases, the constraints become increasingly stringent, scaling as $m_{a}^{-1/2}$. The enhancement of constraints comes from two facts. First, the background effect modifies the scaling of the potential from $1/r^{3}$ to $1/r$. Second, the effect is proportional to the number density of axion DM and is inversely proportional to the axion mass. The dashed extension lines attached to the solid lines represent the parameter region where the DM de Broglie wavelength is comparable with the separation between the test and source masses in the two experiments. As $m_{a}$ increases and thus the de Broglie wavelength becomes correspondingly shorter, the constraints are expected to smoothly revert to the vacuum case.
Discussion and conclusion– In this work, we have investigated the impact of the axion DM background on the spin-independent axion forces between nucleons. This is particularly important for the axion case, as the background effect modifies only the loop-level results, and the spin-independent axion forces themselves are loop-level phenomena. By systematically calculating the relevant Feynman diagrams, we have demonstrated that the axion force is significantly amplified in proportion to the number density of axion DM particles. Furthermore, the scaling of the potential for this force is modified from $1/r^{3}$, as in the vacuum-only case, to $1/r$ for distance $r$ smaller than the DM de Broglie wavelength. Consequently, across a broad range of axion masses, constraints on the axion parameter $c_{GG}/f$ have been significantly enhanced by existing results from fifth-force experiments, including Casimir-less and torsion balance tests. This underscores the fact that these types of experiments are more effective than previously believed in searching for the hypothetical particle, the axion.
An interesting future direction would involve investigating how the background effect diminishes when the typical length scale of experiments approaches or exceeds the DM de Broglie wavelength. Additionally, quantum sensors, such as atomic interferometers Abend:2020djo ; Bass:2023hoi , atomic clocks Peik:2020cwm ; Blaum:2021eay , or emerging experimental proposals  Carney:2021irt ; Budker:2021quh ; Fierlinger:2024rdj , may provide further insights on testing the background force of the axion. It should be noted that for simplicity, our calculations assume an isotropic distribution of DM. A more meticulous analysis may reveal daily and/or annual modulations of the enhanced force.
@@ Acknowledgements
We thank Bingrong Yu for his discussion on the decoherence effect and Kevin Zhou for providing his insight into a possible classical description of the background force after the first version of this paper appears. This work is partly supported by the National Research Foundation of Korea (NRF) under Grant No. RS-2024-00405629.